find the indicated outputs f(x)=4x^2-5x
f(0)=
f(-1)=
f(2)=

Found 2 solutions by checkley77, Alan3354:
Answer by checkley77(12569)

f(x)=4x^2-5x
f(0)=4*0^2-5^0=4*0-5*0=0-0=0
f(-1)=4*-1^2-5*1=4*1*-5=-1
f(2)=4*2^2-5*2=4*4-10=16-10=6

Answer by Alan3354(30993)

f(x)=4x^2-5x
f(0)=
f(-1)=
f(2)=
----------
Sub the values for x
f(0) = 4*0^2 - 5*0
f(0) = 0
----------
f(-1) = 4*(-1)^2 -5*(-1)
= 4 + 5
= 9
-------------
f(2) = 4*2^2 - 5*2
= 16 - 10
= 6